OpenStudy (anonymous):

In a proton linear accelerator, a 8.6 mA proton current hits a target. (a) How many protons hit the target each second? (b) What is the energy delivered to the target each second if the protons each have a kinetic energy of 20 MeV and come to a complete stop in the target? J/s

OpenStudy (anonymous):

An amp is one coulomb (~6.241×10^18 times the elementary charge) per second. A proton has an elementary charge of 1. Use this to find how many protons per second are hitting the target. Once you have that, since the protons are coming to a complete stop, they are each imparting all of their kinetic energy into the target. Use the number of protons per second and the kinetic energy of each to find the energy delivered to the target each second (in MeV). You'll then want to convert MeV to Joules. \[1\, \mathrm{eV} = 1.60217657 × 10^{-19}\, \mathrm{joules}\]

OpenStudy (michele_laino):

Hint: a proton carries 1.6*10^(-19)Couloms of electricity, now we have a current of 8.6*10^(-3) Coulombs/sec, so the requested number of proton is: \[\Large N = \frac{{8.6 \times {{10}^{ - 3}}}}{{1.6 \times {{10}^{ - 19}}}} = ...?\]

OpenStudy (michele_laino):

Now, if each proton has 20 MeV of energy, then the total energy E gained by our target, is: \[\Large E = 20 \times N = ...{\text{MeV}}\]